use the product-to-sum formula to rewrite the product as a sum or difference. sin(9x) sin(4x)
Answer #1

o Sin 93) Sln(431) 2 SIA SinB = cos(6-8) - COS(A+B) = 3 [asmon) seme-x3] = { [cos(9*- 43) - CO3(936+43)] = = cos(5) – COS(.33Thank